# A meaningful prediction of functional decline in amyotrophic lateral sclerosis based on multi-event survival analysis

### v1

@@ Introduction
Discrimination: We want good discrimination performance if we need to rank patients by their disease severity to prioritize access to limited clinical resources, for example, portable ventilators. Across all models (Table 4), we see the highest Harrell’s and Uno’s C-indices for the Speech (approx. 74–75% and 70–71%, respectively) and Swallowing (approx. 75–76% and 70–71%, respectively) events. On the other hand, functional decline with respect to limb functions are harder to discriminate, with lower Harrell’s/Uno’s CI (approx. 67–69%) for Handwriting and Walking events. Although Dyspnea typically has moderate CI values (approx. 66–68%), RSF achieves a Harrell’s CI of 78.23% and a Uno’s CI of 78.95% for this event, making it the best model in this case. DeepSurv and MTLR also demonstrate consistently good discrimination performances across events, and MENSA matches or slightly exceeds these for Speech and Swallowing, though its performance dips slightly for Walking and Dyspnea.
Absolute/squared error: We want accurate time-to-event estimates or survival probabilities when a decision or intervention depend on the precise timing of the event. Across all models (Table 4), the Swallowing event consistently exhibits some of the lowest IBS (approx. 12-13) and mMAE (approx. 110-150) numbers. For Dyspnea, the IBS is also moderate (11.9-13.8), but the RSF is notable with a low IBS along with a relatively high mMAE (approx. 850$\pm$471 days). In contrast, other models produce much more moderate mMAE for this event (approx. 200-400 days), as seen with the MENSA model. This highlights an important point: simply calculating the squared error of the predicted survival probabilities does not fully capture the model’s ability to predict the time to event. MENSA achieves reasonable mMAE (approx. 120–190 days) in all events, except for Speech (181$\pm$8 days), and tends to perform particularly well for Swallowing, similar to MTLR.
@@
Fig. 3 shows the prediction error in days between the population-level KM estimator [30] and the covariate-based models. The KM estimate is calculated per event on observed and unobserved event times, without considering individual differences (i.e., covariates). In all five events, the covariate-based models generally give superior time-to-event estimates compared to the KM method, with some variation between models. This suggests that incorporating patient-specific information can lead to more accurate time-to-event predictions when predicting functional decline than simply relying on a population-level estimate.
@@
Calibration: We want calibrated event probabilities when we have to use them to guide our decision-making. For example, underestimating the risk of respiratory decline or bulbar involvement may result in mistimed interventions or suboptimal treatment choices. Across all models (Table 4), the Walking event consistently poses a calibration challenge, as predicted event probabilities generally deviate from the observed event frequencies (i.e., the models are sometimes not calibrated). However, RSF achieves perfect D-calibration (10/10) across all events and experiments, demonstrating robust and consistent calibration performance. DeepSurv and MTLR also perform well, with most experiments achieving D-calibration. MENSA achieves perfect calibration for Swallowing and Dyspnea (10/10), but shows slightly lower calibration counts for Speech (9/10), Handwriting (9/10), and especially Walking (only 6/10).
{adjustwidth}
-2.25in0in Model Event Harrell’s CI $\uparrow$ Uno’s CI $\uparrow$ IBS $\downarrow$ mMAE $\downarrow$ D-Cal $\uparrow$ CoxPH [21] Speech 74.52$\pm$1.39 70.44$\pm$3.34 14.41$\pm$0.43 206.85$\pm$9.88 (10/10) Swallowing 75.75$\pm$1.53 70.96$\pm$4.58 12.58$\pm$0.72 146.15$\pm$25.50 (10/10) Handwriting 67.57$\pm$0.92 63.75$\pm$2.29 16.73$\pm$0.70 141.85$\pm$7.40 (10/10) Walking 68.67$\pm$1.20 67.15$\pm$1.40 16.60$\pm$0.50 145.04$\pm$4.50 (7/10) Dyspnea 68.54$\pm$1.71 66.99$\pm$6.30 13.79$\pm$0.41 235.46$\pm$27.57 (10/10) RSF [22] Speech 75.05$\pm$1.13 70.08$\pm$2.75 14.69$\pm$0.44 164.51$\pm$19.12 (10/10) Swallowing 75.53$\pm$1.46 69.65$\pm$2.98 13.02$\pm$0.74 111.71$\pm$11.46 (10/10) Handwriting 67.38$\pm$1.39 63.70$\pm$2.05 16.87$\pm$0.57 146.97$\pm$6.88 (10/10) Walking 68.44$\pm$1.47 66.68$\pm$1.54 16.78$\pm$0.56 145.98$\pm$4.63 (10/10) Dyspnea 78.23$\pm$1.07 78.95$\pm$4.01 11.89$\pm$0.33 850.45$\pm$471.39 (10/10) DeepSurv [23] Speech 74.18$\pm$3.05 70.77$\pm$3.57 14.36$\pm$0.89 232.15$\pm$28.48 (10/10) Swallowing 75.61$\pm$1.37 70.67$\pm$5.52 12.48$\pm$0.65 178.77$\pm$39.68 (10/10) Handwriting 67.12$\pm$1.08 63.72$\pm$2.15 16.81$\pm$0.84 151.83$\pm$18.59 (9/10) Walking 69.53$\pm$0.93 68.44$\pm$1.11 16.23$\pm$0.51 148.17$\pm$6.60 (8/10) Dyspnea 71.72$\pm$3.99 70.61$\pm$7.60 13.11$\pm$0.61 384.96$\pm$284.17 (10/10) MTLR [24] Speech 74.22$\pm$1.67 70.17$\pm$3.23 13.95$\pm$0.44 179.35$\pm$5.34 (10/10) Swallowing 75.13$\pm$1.43 70.63$\pm$5.41 12.34$\pm$0.59 122.93$\pm$19.52 (10/10) Handwriting 67.32$\pm$1.11 63.71$\pm$2.42 16.72$\pm$0.66 140.94$\pm$6.29 (9/10) Walking 68.97$\pm$1.16 67.71$\pm$1.06 16.30$\pm$0.51 149.93$\pm$10.99 (9/10) Dyspnea 66.54$\pm$2.22 65.90$\pm$6.85 13.50$\pm$0.50 203.36$\pm$24.57 (10/10) MENSA [25] Speech 74.21$\pm$1.11 70.56$\pm$3.47 14.53$\pm$0.41 181.81$\pm$7.92 (9/10) Swallowing 74.94$\pm$1.36 70.85$\pm$5.20 12.61$\pm$0.56 119.16$\pm$7.12 (10/10) Handwriting 66.46$\pm$0.90 63.80$\pm$2.65 17.10$\pm$0.66 141.97$\pm$5.70 (9/10) Walking 68.63$\pm$1.29 66.94$\pm$1.50 17.04$\pm$0.61 149.50$\pm$6.49 (6/10) Dyspnea 66.56$\pm$1.77 64.40$\pm$7.07 13.78$\pm$0.30 189.43$\pm$24.28 (10/10)
@@
Table 4: Empirical results, averaged over ten independent experiments ($\pm$ SD). D-calibration counts the number of times the respective model was D-calibrated. Harrell’s CI, Uno’s CI and IBS results are multiplied by 100 for better readability.
{adjustwidth}
-2.25in0in Refer to caption
@@
Figure 3: The mMAE (in days) as a function of covariate-based models (x-axis) in the PRO-ACT test set, with error bars representing empirical 95% confidence intervals. The horizontal dashed line is the KM estimator [30]. Lower is better.
@@
Our method can predict an individual’s disease trajectory as $K$ individual survival distributions, where $K$ is the number of events (here, $K=5$). Fig. 4 shows such ISDs, from baseline ($t=0$) to 500 days later, for the $i$-th patient (say Mr. Smith). Mr. Smith is a 72-year-old male, who had been diagnosed with limb-onset ALS prior to study entry and had experienced difficulty speaking and weakness in his legs, but all five of his scores were above 2, so he could participate in the study. Mr. Smith’s covariates are available in the Supplement (S4 Table.). The neurologist predicted that speech and walking difficulties would be among Mr. Smith’s earliest functional impairments, and our method confirmed her suspicion. From the outset, Mr. Smith was provided with a walking aid and a personalized long-term care plan that included therapy, support systems, and home modifications. Shortly after his hospital visit, he had to repeat many of his words to be understood and relied on assistance from a cane or his wife to walk. Meanwhile, handwriting also became challenging for him, but he did not experience shortness of breath or difficulty swallowing whole foods until much later.
Figure 4: Predicted ISDs for Mr. Smith ($i$). The point where the survival curve intersects the dashed horizontal line at 50% indicates the predicted time of event – so this predicts, for example, Walking at approximately 80 days, Handwriting at 240 days, and Speech at 300 days. The dashed vertical lines are Mr. Smith’s actual time of event for the respective events – so Speech at approximately 7 days, Handwriting at 35 days, and Walking at 45 days.
@@
Our method can also make individual counterfactual predictions [32], that is, alternative scenarios, in which certain patient covariates are changed to see their (predicted) effect on the probability of the event. Such predictions can give novel insights into disease mechanisms and assess treatment options on an individual level. In the following figures, we have used MENSA [25] to make individual counterfactual predictions based on the data in the PRO-ACT test set. These provide insights into the relationship between common disease predictors and functional decline, and asks the question how changing certain covariates can affect disease outcome. Again, we use our fictional patient, Mr. Smith, as the instance $i$. Additionally, we separate all patients in the test set into groups by the same covariate we used for the individual patient and use the logrank test [33] to measure the statistical difference between the groups’ event times. We use a Bonferroni correction to correct for multiple comparisons by dividing the significance level ($\alpha=0.05$) by the number of comparisons ($n=5$), i.e., the significance level is then $\alpha_{adjusted}=0.05/5=0.01$.
@@
Many newly-diagnosed patients with ALS need to decide whether to take certain pharmaceutical drugs. Currently, there are three approved drugs by the US Food and Drug Administration (FDA) to treat ALS, the most common being Riluzole. Riluzole works by blocking the release of glutamate, which delays the onset of ventilator-dependence or tracheostomy in some people and may increase overall survival by two to three months [34]. However, the drug can have significant side effects, such as nausea, anorexia, and diarrhea [4], which makes it relevant to determine whether the drug will be effective -– i.e., significantly delay functional decline – for each person, to determine whether that person should take this drug. Upon diagnosis, Mr. Smith decided not to take Riluzole, but what if he, instead, had decided to take Riluzole. Fig. 5 shows the predicted individual survival distributions (ISDs) for each of the five events using Mr. Smith’s covariates ($\bm{x}_{i}$), with the Riluzole covariate being yes (resp., no) at the baseline visit. The two survival curves have nearly the same slopes, with Riluzole only being marginally above non-Riluzole, and nearly identical estimated event times. There are also no statistical differences in the event times between the groups either ($p>0.01$). As an example, for the Speech event, this model predicts that taking Riluzole would delay the onset of speaking difficulties by approximately 20 days.
@@
In ALS, the site of onset refers to the specific part of the body where the initial symptoms began. ALS can affect different regions and the site of onset can influence the progression of the disease. The two most common sites of onset are bulbar-onset and limb-onset, the former being associated with faster disease progression in general [11]. Mr. Smith was diagnosed with limb-onset ALS, but what if instead he had been diagnosed with bulbar-onset ALS. Fig. 6 shows the predicted ISDs for Mr. Smith based on whether he has limb or bulbar-onset. We see significant differences in the prediction for the Speech and Swallowing events, indicating a worse prognosis for Mr. Smith if he has bulbar-onset, but more agreement between the curves for the Handwriting, Walking and Dyspnea events.
@@
Forced Vital Capacity (FVC) is a measure of lung function that represents the total amount of air a person can forcibly exhale after taking a deep breath. It is commonly used to assess lung respiratory function and is measured in liters. In ALS, FVC is a key clinical measure for evaluating respiratory muscle strength, which progressively weakens as the disease progresses. Mr. Smith’s total FVC score at baseline was 2.69. Again, we ask the question of what if instead he had a higher FVC or a lower FVC. Fig. 7 shows the predicted ISDs for Mr. Smith based on whether he has a high or low FVC. As expected, the Handwriting event shows little significant relationship with FVC, as evidenced by the high $p$-value and the alignment of the two survival curves. In contrast, the Swallowing event exhibits a significant relationship with FVC, and the model provides widely different time-to-event predictions based solely on this covariate.
@@
Lastly, we consider the total ALSFRS-R score at baseline and its effect on predicted functional decline. This score provides a snapshot of Mr. Smith’s initial physical capabilities when he is relatively healthy at before the disease has advanced notably. Mr. Smith’s total ALSFRS-R score at baseline was 37. Again, we make a counterfactual prediction if instead Mr. Smith had a significantly higher or lower ALSFRS-R score. Fig. 8 shows the predicted ISDs for Mr. Smith based on whether he has a high or low ALSFRS-R score. Consistent with the literature [4], if Mr. Smith had a high ALSFRS-R score to begin with, he would (presumably) experience functional decline much later than if he had a low ALSFRS-R score. This is indicated by the discrepancy between the survival curves and the low $p$-values across all the events. Obviously, initial physical ability has a close relationship with loss of function after a certain time, and patients may be at different stages of the disease and have different levels of strength when they enter the trial. In addition, research in ALS has suggested the presence of various disease phenotypes with different progression rates [35].
{adjustwidth}
-2.25in0in Refer to caption
@@ Discussion
This research has three goals: (1) determine whether covariate-based survival models can predict the time to individual functional decline (ALSFRS-R $\leq$ 2) in ALS patients more accurately than population estimators (e.g., Kaplan-Meier), (2) precisely estimate risk scores and event times, and (3) explore how counterfactual scenarios (e.g., what if the patient took Riluzole) affect the probability of functional decline over the course of the disease.
@@
For the first goal, individualized predictions are crucial for tailoring healthcare and therapeutic decisions to the individual patient. Models that estimate an individual’s risk rather than a group average enable us to answer questions such as: how long will I be able to speak clearly enough to be understood? We make a novel contribution in this regard, as we empirically show that most covariate-based models, even simple linear models, outperform the Kaplan-Meier estimator (which does not use covariates) in predicting the time to functional decline. Given that ALS affects multiple functional systems (bulbar, motor, respiratory) at varying rates, our analysis also shows clear advantages of adopting non-proportional hazards models, such as Random Survival Forests and MENSA. Unlike state-of-the-art methods for predicting functional decline in ALS [12, 9, 10, 11, 3], our method is unique in that it provides patients and clinicians with an actual time (in days) when functional decline is likely to occur.
@@
For the second goal, we rigorously evaluate five survival models across three aspects of model performance: discrimination, absolute/squared error, and calibration. The main takeaway is that some types of functional decline in ALS patients are easier to predict than others. For example, the Speech event has an average Harrell’s CI of approximately 74% across five models – the highest among all events – but also an average mMAE (using margin times for censored instances) of 192.4 days, the second-highest. This means that even if a model is good at ranking patients by relative risk (independent of time) for some event, it cannot necessarily predict when the event will occur. In contrast, the Walking event has a lower average CI of approximately 67%, but an average mMAE of only 147.7 days. This suggests that, although it is harder to rank patients by risk for this event compared to Speech, it is easier, by contrast, to predict when the event will occur. This leads to a novel observation that strong ranking performance does not necessarily translate into accurate time-to-event predictions in this case.
@@
Finally, for the third goal, we use our method to make counterfactual predictions, investigating how modifying specific covariates affects disease outcomes for individual patients. To our knowledge, this is a novel contribution to ALS research, as no prior work has explored this use case. However, such predictions are highly relevant for medical professionals who must make treatment decisions after first meeting a patient. For example, consider our hypothetical patient, Mr. Smith: based on his covariates, there were no significant differences in the predicted survival (event) probabilities should he decide to take Riluzole, a drug used to treat ALS. In other words, whether he decided to take Riluzole or not, this decision would not significantly change his predicted risk of functional decline over the next 500 days from the first time he saw the neurologist. We believe that these individualized predictions can be very beneficial for treatment planning and clinical decision-making.
@@
Our method has several important applications. First, it can help medical doctors make informed decisions by tailoring treatments and interventions to the individual’s predicted trajectory. This is crucial for ALS, where timely interventions, such as assistive devices and nutritional or respiratory support, can significantly impact a patient’s quality of life. Second, patients and their families are often faced with difficult decisions regarding treatment options, lifestyle adjustments, and end-of-life planning. By providing a clear, individualized prognosis, clinicians can help patients make more informed decisions about their care, improve their quality of life, and give them a sense of control over their treatment. As part of the evaluation, we show how our method can make counterfactual predictions, answering the what-if questions: Mr. Smith was diagnosed with limb-onset ALS and the model predicts a specific path of functional decline for him, but what would have happened if he had been diagnosed with bulbar-onset ALS instead. Third, personalized predictions can also help design and evaluate clinical trials. Understanding the likely timing of functional decline in patients with ALS can help researchers identify suitable candidates for clinical studies, ensuring that the trials are conducted with a population in the appropriate stage of the disease. This can lead to more effective testing of potential therapies and interventions that could slow or stop disease progression.
@@
    Integrate model training with existing electronic health record (EHR) systems or clinical decision support tools to automatically retrieve patient data (covariates) and past ALSFRS-R scores (event information). After obtaining the relevant data, they should train the model, evaluate it, deploy it, and expose an application programming interface (API) that allow clinicians to make predictions on new patients.
@@
    Develop a user-friendly software interface or web-based application that allows clinicians to easily input or access patient information and obtain individualized predictions of functional decline over time. When a new patient is diagnosed with ALS and has their first visit at the hospital, clinicians can use the interface to enter the patient’s covariates and obtain a prediction over the next 500 days.
@@
    Update, retrain and deploy the model regularly with new patient data to maintain accuracy and incorporate new biomarkers or additional clinical variables, if such become available. This can be done using Machine Learning Operations (MLOps), which are a set of practices that automate and simplify machine learning workflows and deployments.
@@
All survival models evaluated in this work assume (conditional) independent censoring, i.e., the event time is independent of the censoring time, given the patient’s covariates. For example, patients who leave the study early often do so for reasons unrelated to the event of interest, or because of poor lung function captured by the forced vital capacity covariate. However, ALS is a heterogeneous disease that can quickly deteriorate and no study can capture all intrinsic disease mechanisms in the recorded covariates. Generally, clinical trials in ALS experience high dropout rates [36], which is also evident in our study through the high censoring rates shown in Table 3 and Fig. 2. It is likely that this phenomenon is related to disease progression and hereby functional decline, but since we do not observe functional decline in patients who leave the study, this kind of dependent censoring may bias the results. According to Atassi et al. [36], the three most common reasons for attrition in ALS trials were withdrawal of consent for no specific reason (37%), death (28%) and withdrawal secondary to an adverse event (17%).
@@
Our method has several avenues for future research. First, addressing the aforementioned dependence between the event and censoring distributions is relevant, given the relatively few covariates provided in the PRO-ACT dataset and the fact that all of the evaluated survival models assume that censoring is conditionally independent of the event given the covariates. To address this issue, when estimating model parameters, we can also learn a so-called copula [37]. A copula is a link function that describes the dependence structure between the event and censoring distributions. Copulas have recently gained interest in survival analysis due to their flexibility in capturing complex dependence structures [38]. Second, it is highly relevant to include covariates captured using magnetic resonance imaging (MRI) into the predictive models, since such features can explain early signs of neurodegeneration, for example, cerebral atrophy in the motor cortex and corticospinal tract [39]. At present, such features have only been used to predict mortality or disease aggressiveness in patients with ALS [39, 19, 40], not functional decline. Thus, this presents an exciting avenue for future research.

### v2

@@ Introduction
Discrimination: Good discrimination performance is important if we need to rank patients by their disease severity to prioritize access to limited clinical resources, for example, portable ventilators. Across all models, we see the highest Harrell’s and Uno’s C-indices for the Speech (approx. 74–75% and 70–71%, respectively) and Swallowing (approx. 75–76% and 70–71%, respectively) events. On the other hand, functional decline with respect to limb functions is harder to discriminate, with lower C-indices (approx. 67–69%) for the Handwriting and Walking events. Although the Dyspnea event typically has moderate CI values (approx. 66–68%), RSF achieves a Harrell’s CI of 78.23% and a Uno’s CI of 78.95% for this event, making it the best model in this case. DeepSurv and MTLR also demonstrate consistently good discrimination performance across events, and MENSA matches or slightly exceeds these for the Speech and Swallowing events, though its performance dips slightly for the Walking and Dyspnea events.
[…]
@@
Fig. 3 shows the prediction error in days between the population-level KM estimator [30] and the covariate-based models. The KM estimate is calculated per event on observed and censored event times, without considering individual differences (i.e., covariates). In all five events, the covariate-based models generally give superior time-to-event estimates compared to the KM method, with some variation between models. This suggests that incorporating patient-specific information can lead to more accurate time-to-event predictions when predicting functional decline, rather than simply relying on a population-level estimate.
@@
Calibration: Good calibration performance is important when we have to use the actual predicted probabilities to guide our decision-making. For example, underestimating the risk of respiratory decline or bulbar involvement may result in mistimed interventions or suboptimal treatment choices. Across all models, the Walking event consistently poses a calibration challenge, as predicted event probabilities generally deviate from the observed event frequencies here (i.e., the models are sometimes not calibrated). However, RSF achieves perfect D-calibration (10/10) across all events and experiments, demonstrating robust and consistent calibration performance. DeepSurv and MTLR also perform well, with most experiments achieving D-calibration. MENSA achieves perfect calibration for the Swallowing and Dyspnea (10/10) events, but shows slightly poorer calibration for the Speech (9/10) and Handwriting (9/10) events, and especially the Walking event (only 6/10).
@@
Table 4: Mean ($\pm$ SD.) prediction performance, averaged over 10 independent experiments. D-calibration counts the number of times the respective model was D-calibrated. Harrell’s CI, Uno’s CI and IBS results are multiplied by 100.
Model
Event
Harrell’s CI $\uparrow$
Uno’s CI $\uparrow$
IBS $\downarrow$
mMAE $\downarrow$
D-Cal $\uparrow$
CoxPH
[21]
[…]
RSF
[22]
[…]
DeepSurv
[23]
[…]
MTLR
[24]
[…]
MENSA
[25]
[…]
[Fig. 3]
@@
Figure 3: The mMAE (in days) as a function of covariate-based models (x-axis) in the PRO-ACT test set, with error bars representing empirical 95% confidence intervals. The horizontal dashed line is the KM estimator. Lower is better.
@@
Our method can predict a patient’s disease trajectory as $K$ individual survival distributions, where $K$ is the number of events (here, $K=5$). Fig. 4 shows such ISDs, from baseline ($t=0$) to 500 days later, for the $i$-th patient. As an example, let us return to our patient, Mr. Smith. He is a 72-year-old male, who had been diagnosed with limb-onset ALS prior to study entry and had experienced difficulty speaking and weakness in his legs, but all five of his scores were above 2, so that he could participate in the study. Mr. Smith’s covariates are available in the Supplement (S4 Table.). The neurologist predicted that talking and walking difficulties would be among Mr. Smith’s earliest functional impairments, and our method confirmed her suspicion. From the outset, Mr. Smith was provided with a walking aid and a personalized long-term care plan that included therapy, support systems, and home modifications. Shortly after his hospital visit, he had to repeat many of his words to be understood and relied on assistance from a cane or his wife to walk. Meanwhile, handwriting also became challenging for him, but he did not experience shortness of breath or difficulty swallowing whole foods until much later.
[Fig. 4]
Figure 4: Predicted ISDs for Mr. Smith (ii). The point where the survival curve intersects the dashed horizontal line at 50% indicates the predicted time of event – so this predicts, for example, Walking at approximately 80 days, Handwriting at 240 days, and Speech at 300 days. The dashed vertical lines are Mr. Smith’s actual time of event for the respective events – so, for example, Speech at approximately 7 days, Handwriting at 35 days, and Walking at 45 days.
@@
Our method can also make individual counterfactual predictions [32], that is, alternative scenarios in which certain patient covariates are changed to see their (predicted) effect on the probability of the event. Such predictions can give novel insights into disease mechanisms and assess treatment options on an individual level. In the following figures, we have used MENSA [25] to make individual counterfactual predictions based on our PRO-ACT dataset. These provide insights into the relationship between common disease predictors and functional decline, and ask how changing certain covariates affects disease outcomes. Again, we use our patient, Mr. Smith, as the instance $i$. Additionally, we separate all patients in the test set into groups by the same covariate we used for the individual patient and employ the log-rank test [33] to measure the statistical difference between the groups’ event times. We use a Bonferroni correction to correct for multiple comparisons by dividing the significance level ($\alpha=0.05$) by the number of comparisons ($n=5$), i.e., the significance level is then $\alpha_{adjusted}=0.05/5=0.01$. For the remainder of this section, we will discuss our counterfactual predictions.
@@
As a starting point, many newly diagnosed patients with ALS need to decide whether to take certain pharmaceutical drugs. Currently, there are three approved drugs by the US Food and Drug Administration (FDA) to treat ALS, the most common being Riluzole. Riluzole works by blocking the release of glutamate, which delays the onset of ventilator dependence or tracheostomy in some people and may increase overall survival by two to three months [34]. However, the drug can cause significant side effects, including nausea, anorexia, and diarrhea [4]. Therefore, it is important to determine whether the drug will be effective – i.e., significantly delay functional decline in this regard – for each individual, in order to decide whether that person should receive the treatment. Upon diagnosis, Mr. Smith decided not to take Riluzole, but what if he instead had decided to take it. Fig. 5 shows the predicted ISDs for each of the five events using Mr. Smith’s covariates ($\bm{x}_{i}$), with the Riluzole covariate being yes (resp., no) at the baseline visit. The two survival curves have nearly the same slopes, with Riluzole only being marginally above non-Riluzole, and nearly identical estimated event times. There are no statistical differences in event times between the groups either ($p>0.01$). As an example, for the Speech event, this model predicts that taking Riluzole would delay the onset of speaking difficulties by approximately 20 days.
@@
In ALS, the site of onset refers to the specific part of the body where the initial symptoms began. ALS can affect different regions, and the site of onset can influence the progression of the disease. The two most common sites of onset are bulbar-onset and limb-onset, the former being associated with faster disease progression in general [11]. Mr. Smith was diagnosed with limb-onset ALS, but what if instead he had been diagnosed with bulbar-onset ALS. Fig. 6 shows the predicted ISDs for Mr. Smith based on whether he has limb-onset or bulbar-onset. We see significant differences in the predictions for the Speech and Swallowing events, indicating a worse prognosis for Mr. Smith if he has bulbar-onset, but more agreement between the curves for the Handwriting, Walking and Dyspnea events.
@@
Next, forced Vital Capacity (FVC) is a measure of lung function and represents the total amount of air a person can forcibly exhale after taking a deep breath. It is commonly used to assess lung respiratory function and is measured in liters. In ALS, FVC is a key clinical assessment for evaluating respiratory muscle strength, which progressively weakens as the disease progresses. Mr. Smith’s total FVC score at baseline was 2.69. Again, we ask the counterfactual question of what if instead he had a higher FVC or a lower FVC. Fig. 7 shows the predicted ISDs for Mr. Smith based on whether he has a high or low FVC. As expected, the Handwriting event shows only a weak relationship with FVC, as evidenced by the high $p$-value and the alignment of the two survival curves. In contrast, the Swallowing event exhibits a significant relationship with FVC, and the model provides substantially different time-to-event predictions based solely on this covariate.
@@
Lastly, we consider the total ALSFRS-R score at baseline and its effect on predicted functional decline. This score provides a snapshot of Mr. Smith’s initial physical capabilities when he is relatively healthy before the disease has advanced notably. Mr. Smith’s total ALSFRS-R score at baseline was 37. Again, we make a counterfactual prediction if instead Mr. Smith had a significantly higher or lower ALSFRS-R score. Fig. 8 shows the predicted ISDs for Mr. Smith based on whether he has a high or low ALSFRS-R score. Consistent with the literature [4], if Mr. Smith had a high ALSFRS-R score to begin with, he would (presumably) experience functional decline much later than if he had a low ALSFRS-R score. This is indicated by the discrepancy between the survival curves and the low $p$-values across all the events. Obviously, initial physical ability has a strong relationship with loss of function after a certain time, and patients may be at different stages of the disease and have different levels of strength when they enter the trial. In addition, research in ALS has suggested the presence of various disease phenotypes with different progression rates [35].
[Fig. 5]
@@ Discussion
This research has three goals: (1) determine whether covariate-based survival models can predict the time to individual functional decline (ALSFRS-R $\leq$ 2) in ALS patients more accurately than population estimators, (2) precisely estimate risk scores and event times, and (3) explore how counterfactual scenarios affect the probability of functional decline over the course of the disease.
@@
For the first goal, individualized predictions are crucial for tailoring healthcare and therapeutic decisions to the individual patient. Models that estimate an individual’s risk rather than a group average enable us to answer important questions from the patient’s point of view such as: how much longer will I be able to speak clearly enough to be understood? We make a novel contribution in this regard, as we empirically show that most covariate-based models, even simple linear models, outperform the KM estimator (which does not use covariates) in predicting the time to functional decline. Given that ALS affects multiple functional systems (bulbar, motor, respiratory) at varying rates, our analysis also shows clear advantages of adopting non-proportional hazards models, such as Random Survival Forests and MENSA. Unlike other methods for predicting functional decline in ALS [12, 9, 10, 11, 3], our method uniquely provides patients and clinicians with an actual time (in days) when functional decline is likely to occur.
@@
For the second goal, we rigorously evaluate five survival models across three aspects of model performance: discrimination, absolute/squared error, and calibration. The main takeaway is that some types of functional decline in ALS patients are easier to predict than others. For example, the Speech event has an average Harrell’s CI of approximately 74% across the five models – the highest among all events – but also an average mMAE (a variant of the MAE that supports censoring) of 192.4 days, the second-highest. This means that even if a model is good at ranking patients by their relative risk (independent of time) for some event, it cannot necessarily predict accurately when the event will occur. In contrast, the Walking event has a lower average CI of approximately 67%, but an average mMAE of only 147.7 days. This suggests that, although it is harder to rank patients by risk for this event compared to Speech, it is easier, by contrast, to predict when the event will occur. This leads to the realization that strong ranking performance does not necessarily translate into accurate time-to-event predictions.
@@
Finally, for the third goal, we use our method to make counterfactual predictions, investigating how modifying specific covariates affects predicted disease outcomes for individual patients. To our knowledge, no prior work has explored this use case in ALS. However, such predictions are highly relevant for medical professionals who must make important treatment decisions after first meeting a patient. For example, consider our patient, Mr. Smith: based on his covariates, there were no significant differences in the predicted survival probabilities (i.e., the risk of functional decline) should he decide to take Riluzole, a drug used to treat ALS. In other words, whether he decided to take Riluzole or not, this decision would not significantly change his predicted risk of functional decline over the next 500 days from the first time he saw the neurologist. We believe that these individualized predictions can be very beneficial for treatment planning and clinical decision-making.
@@
Our method has several important applications. First, it can help medical doctors make informed decisions by tailoring treatments and interventions to the individual’s predicted disease trajectory. This is crucial in ALS, where timely interventions, such as assistive devices and nutritional or respiratory support, can significantly impact a patient’s quality of life. Second, patients and their families are often faced with difficult decisions regarding treatment options, lifestyle adjustments, and end-of-life planning. By providing a clear, individualized prognosis, clinicians can help patients make more informed decisions about their care, improve their quality of life, and give them a sense of control over their treatment. As part of the evaluation, we show how our method can make counterfactual predictions, answering the what-if questions: Mr. Smith was diagnosed with limb-onset ALS and the model predicts a specific path of functional decline for him, but what would have happened if he had been diagnosed with bulbar-onset ALS instead. Third, personalized predictions can also help design and evaluate clinical trials. Understanding the likely timing of functional decline in patients with ALS can help researchers identify suitable candidates for clinical studies, ensuring that the trials are conducted with participants at the appropriate stage of disease. This can lead to more effective testing of potential therapies and interventions that could slow or stop disease progression.
@@
    Integrate model training with existing electronic health record (EHR) systems or clinical decision support tools to automatically retrieve patient data (covariates) and past ALSFRS-R scores (event information). After obtaining the relevant data, they should train the model, evaluate it, deploy it, and expose an application programming interface (API) that allows clinicians to make predictions on new patients.
@@
    Develop a user-friendly software interface or web-based application that allows clinicians to easily input or access patient information and obtain individual predictions of functional decline over time. When a new patient is diagnosed with ALS and has their first visit at the hospital, clinicians can use the application to enter the patient’s covariates and obtain a prediction over the next 500 days.
@@
    Update, retrain, and deploy the model regularly with new patient data to maintain accuracy and incorporate new biomarkers or additional clinical variables, if such become available. This can be done using Machine Learning Operations (MLOps), which is a set of practices that automate and simplify machine learning workflows and deployments.
@@
All survival models evaluated in this work assume conditional independent censoring, i.e., that the event time is independent of the censoring time, given the patient’s covariates. For example, patients who leave the study early do so for reasons unrelated to the event of interest or because of poor lung function captured by the forced vital capacity covariate. However, ALS is a heterogeneous disease that can quickly deteriorate and no study can capture all intrinsic disease mechanisms. Generally, clinical trials in ALS experience high dropout rates [36], which is also evident in our study through the high censoring rates shown in Table 3 and Fig. 2. It is likely that this phenomenon is related to disease progression and therefore functional decline, but since we do not observe functional decline in patients who leave the study, this kind of censoring may violate the assumption of conditional independent censoring and bias the results. According to Atassi et al. [36], the three most common reasons for attrition (i.e., dropout) in ALS trials were withdrawal of consent for no specific reason (37%), death (28%), and withdrawal secondary to an adverse event (17%).
@@
Our method has several avenues for future research. First, addressing the aforementioned dependency between the event and censoring times is relevant, given the relatively few covariates provided in the PRO-ACT dataset and because all evaluated survival models assume that censoring is conditionally independent. To address this issue, when estimating model parameters, we can learn a so-called copula function [37]. This is a link function that describes the dependency structure between the event and censoring distributions. Copulas have recently gained attention in survival analysis due to their flexibility in capturing complex dependency structures [38]. Second, it is highly relevant to include covariates captured using magnetic resonance imaging (MRI) into the predictive models, since such features can explain early signs of neurodegeneration, for example, cerebral atrophy in the motor cortex and corticospinal tract [39]. At present, such features have only been used to predict mortality or disease aggressiveness in patients with ALS [39, 19, 40], but not to predict functional decline. Thus, this presents an exciting avenue for future research.